### v1

# On Springer correspondence for wreath products

Abstract.
We first show that the wreath product $\Sigma_{m}\wr\Sigma_{d}$ between two symmetric groups appears as the generalized Weyl group of an Iwahori’s generalized Tits system. We then introduce a certain subvariety of the flag variety of type A, and then give a geometric proof of its Bruhat decomposition indexed by $\Sigma_{m}\wr\Sigma_{d}$, via the Bialynicki-Birula decomposition. Furthermore, we realize the group algebra $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]$ as the top Borel-Moore homology of a Steinberg variety. Such a geometric realization leads to a Springer correspondence for the irreducible representations over $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$, which can be regarded as a counterpart of the Clifford theory for wreath products.
1. Introduction
1.1. Wreath Products between Symmetric Groups
The quantum wreath products introduced in [LNX23], roughly speaking, are certain quantizations of the group algebras of the wreath product $G\wr\Sigma_{d}$ of a (possibly infinite) group $G$ by the symmetric group $\Sigma_{d}$ on $d$ letters. The one example we want to highlight here is a deformation $\mathcal{H}_{q}({m\wr 2})$ of $\mathbb{C}[\Sigma_{m}\wr\Sigma_{2}]$ which we call the Hu algebra (referred as $\mathcal{A}(m)$ in [Hu02]). The Hu algebra first appeared in Jun Hu’s Morita equivalence theorem between the Hecke algebras of type A and of type $D_{2m}$. It is shown in [LNX23] that $\mathcal{H}_{q}({m\wr 2})$ (and thus its generalizations $\mathcal{H}_{q}({m\wr d})$) shared many favorable properties with the Hecke algebras of type A. In particular, a positivity pattern for the Hu algebra in terms of the dual Kazhdan-Lusztig basis is also observed. It is thus tempting to claim that $\mathcal{H}_{q}({m\wr d})$ should be regarded as the Hecke algebra for the wreath product $\Sigma_{m}\wr\Sigma_{d}$. In this paper, we establish the very first step towards a geometric representation theory for $\Sigma_{m}\wr\Sigma_{d}$.
While the wreath product $\Sigma_{2}\wr\Sigma_{d}$ (i.e., the Weyl group of type B) affords a theory of Hecke algebras, as a special case of the Ariki-Koike algebras [AK94] that quantize the wreath product $C_{m}\wr\Sigma_{d}$; our wreath product $\Sigma_{m}\wr\Sigma_{d}$ between symmetric groups are not complex reflection groups in general, and hence the most general theory of Hecke algebras due to Broué-Malle-Rouquier [BMR98] does not apply.
On one hand, it is well-known that the Ariki-Koike algebras are cyclotomic quotients of the affine Hecke algebras of type A. It is standard (see [CG97]) that the affine Hecke algebras of type A can be obtained via the equivariant K-theory on the Steinberg varities. On the other hand, a classical (i.e., $q\mapsto 1$) result is also available by considering the Borel-Moore homology.
1.2. Bruhat Decompositions for Wreath Products
In order to initiate the geometric representation theory for $\Sigma_{m}\wr\Sigma_{d}$, one needs to construct an algebraic group $G$ admitting a Bruhat decomposition $G=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}BwB$, indexed by $\Sigma_{m}\wr\Sigma_{d}$, for some subgroup $B\subseteq G$.
Our first result is that such a decomposition exists, and we prove it by utilizing Iwahori’s generalized Tits system [I65], which was originally developed to deal with the subtle difference between the extended and non-extended affine Weyl groups (and hence extended versus non-extended Hecke algebras).
1.3. Wreath Flags And Wreath Steinberg Varieties
In the (classical) Tits system $(G,B,N,W)$, the quotient $G/B$ is the well-known flag variety, on which one can develop deep geometric representation theory. In light of our new geneneralized Tits system $(\mathrm{GL}_{m}\wr\Sigma_{d},B_{m}^{d},N_{m\wr d},\Sigma_{m}\wr\Sigma_{d})$, we introduce what we call the wreath flag varieties (see Definition 3.2.1), denoted by $\mathcal{F}l_{m\wr d}$. It carries an action of the wreath product $\mathrm{GL}_{m}\wr\Sigma_{d}$ of the general linear group by a symmetric group, and can be identified with the natural quotient $(\mathrm{GL}_{m}\wr\Sigma_{d})/B^{d}_{m}$.
While one main feature for the classical flag variety with $G=\mathrm{GL}_{m}$ and $B=B_{m}$ is the natural bijection $\mathrm{GL}_{n}/B_{m}\to\mathfrak{B}_{m}\to\mathcal{F}l_{m}$,where $\mathfrak{B}_{m}$ is the variety of all Borel subalgebras in $\mathfrak{gl}_{m}$, and $\mathcal{F}l_{m}$ is the variety of complete flags in $\mathbb{C}^{m}$. We remark that such bijections only partially generalize in our setup for wreath flag variety. The reason is that the obvious analog $\mathfrak{B}_{m\wr d}:=\mathfrak{B}_{m}^{d}$ is not in bijection with the other varieties.
Thus, we have to avoid using arguments involving Borel subalgebras. We then construct a desirable Steinberg variety $Z_{m\wr d}$ via an unconventional Springer resolution. In Table 1 below, we provide a comparison of the two setups.
Type A Construction Wreath Construction
(generalized) Weyl group $\Sigma_{m}$ $\Sigma_{m}\wr\Sigma_{d}$
Lie group $\mathrm{GL}_{m}$ ${G}_{m\wr d}\subseteq\mathrm{GL}_{md}$ Def. 2.2.3
$\cong\mathrm{GL}_{m}\wr\Sigma_{d}$ Prop. 2.4.1
BN-pair $B_{m},~{}N_{m}\subseteq\mathrm{GL}_{m}$ $B_{m}^{d},~{}N_{m}\wr\Sigma_{d}\subseteq G_{m\wr d}$ (2.2.2)
Flag variety $\mathcal{F}l_{m}$ $\mathcal{F}l_{m\wr d}\subseteq\mathcal{F}l_{md}$ Def. 3.2.1
$\cong\mathcal{F}l_{m}\wr\Sigma_{d}$ Prop. 3.2.2(a)
   Homogeneous space $\cong\mathrm{GL}_{m}/B_{m}$ $\cong{G}_{m\wr d}/B_{m}^{d}$ Prop. 3.2.2(b)
   Set of all Borels $\cong\mathfrak{B}_{m}$ irrelevant non-isomorphic
Nilpotent cone $\mathcal{N}_{m}\subseteq\mathfrak{gl}_{m}\curvearrowleft\mathrm{GL}_{m}$ $\mathcal{N}_{m}^{d}\subseteq\mathfrak{gl}_{m}^{d}\curvearrowleft G_{m\wr d}$ (4.1.1)
[…]
Springer resolution $\widetilde{\mathcal{N}}_{m}:=T^{*}\mathcal{F}l_{m}$ $\widetilde{\mathcal{N}}_{m\wr d}:=T^{*}\mathcal{F}l_{m\wr d}$ (4.2.2)
Steinberg variety $Z_{m}:=\widetilde{\mathcal{N}}_{m}\times_{\mathcal{N}_{m}}\widetilde{\mathcal{N}}_{m}$ $Z_{m\wr d}:=\widetilde{\mathcal{N}}_{m\wr d}\times_{\mathcal{N}_{m}^{d}}\widetilde{\mathcal{N}}_{m\wr d}$ Def. 4.3.1
Table 1. A comparison of objects in the two constructions
1.4. Main results
With the above setup, we are able to prove the first few steps towards a geometric representation theory for wreath products, which are summarized below:
Main Theorem.
  1. (a)
    (Corollary 2.3.4, Theorem 3.3.3) There is a Bruhat decomposition of $G$ with Bruhat cells indexed by $\Sigma_{m}\wr\Sigma_{d}$, for some algebraic groups $B,G$. Moreover, it affords a geometric variant $G/B=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}BwB/B$ via the Bialynicki-Birula decomposition.
  2. (b)
    (Theorem 5.2.1) The top Borel-Moore homology of $Z_{m\wr d}$ realizes the group algebra $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]$.
  3. (c)
    (Theorem 6.2.3, Corollary 6.4.3) There is a Springer correspondence between irreducibles over $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$ and certain isotypic components of the top Borel-Moore homology of Springer fibers. Moreover, a connection with the Clifford theory is obtained.
We plan to pursue the (equivariant) K-theory of our Steinberg variety $Z_{m\wr d}$ and its connection with the Hu algebras in a sequel.
2. The Wreath Product $\Sigma_{m}\wr\Sigma_{d}$
2.1. Wreath Products
Denote by $\Sigma_{d}$ the symmetric group on $d$ letters with simple transpositions $t_{j}:=(j~{}j+1)$ for $1\leq j<d$. For any set $X$, we define a set
$X\wr\Sigma_{d}:=X^{d}\times\Sigma_{d}.$ (2.1.1)
For $w\in\Sigma_{d}$, $g_{i}\in X$, by a slight abuse of notation, we use the shorthand notation below:
$(g_{1},\dots,g_{d},1_{\Sigma_{d}})=:(g_{i})_{i},\quad(1_{G},\dots,1_{G},w)=:w\in X\wr\Sigma_{d},$ (2.1.2)
where the later shorthand only makes sense when $X=G$ is a group. In this case, the wreath product carries a group structure $G\wr\Sigma_{d}=G^{d}\rtimes\Sigma_{d}$ from a semidirect product, whose multiplication is determined by
$w(g_{1},\dots,g_{d})=(g_{w^{-1}(1)},\dots,g_{w^{-1}(d)})w\in G\wr\Sigma_{d}\quad\textup{for}\quad w\in\Sigma_{d},g_{i}\in G.$ (2.1.3)
For $g\in G$, we write
$g^{(j)}=(\underset{j-1}{\underbrace{1_{G},\cdots,1_{G}}},g,\underset{d-j}{\underbrace{1_{G},\cdots,1_{G}}})\in G^{d}.$ (2.1.4)
Thus, for $g\in G$, we have $g^{(j+1)}=t_{j}g^{(j)}t_{j}$ in $G\wr\Sigma_{d}$.
Example 2.1.1.
In particular, $\Sigma_{m}\wr\Sigma_{d}$ can be thought as a subgroup of $\Sigma_{md}$ generated by simple transpositions $s_{i}\equiv(i,i+1)$ for $11\leq i\leq m-$ 1 and the following elements $t_{1},\dots t_{m-1}$ of length $m^{2}$:
$\begin{split}t_{j+1}&\equiv\begin{array}[]{|cccccc|}jm+1&\cdots&(j+1)m&(j+1)m+1&\cdots&(j+2)m\\ (j+1)m+1&\cdots&(j+2)m&jm+1&\cdots&(j+1)m\end{array}\\ &=(s_{jm+m}s_{jm+m+1}\cdots s_{jm+2m-1})\cdots(s_{jm+1}s_{jm+2}\cdots s_{jm+m}).\end{split}$ (2.1.5)
Moreover, for $1\leq i,j<d,s\in\Sigma_{m}$,
$t_{i}s^{(j)}t_{i}^{-1}=\begin{cases}s^{(j+1)}&\textup{if }i=j;\\ s^{(j-1)}&\textup{if }i=j-1;\\ s^{(j)}&\textup{otherwise}.\end{cases}$ (2.1.6)
In general, $\Sigma_{m}\wr\Sigma_{d}$ is not a Coxeter group except that $\Sigma_{m}\wr\Sigma_{d}$ is of type $G(m,1,d)$ when $m\leq 2$.
2.2. Tits Systems
Our first step is to show that there is a Bruhat decomposition for some group $G_{m\wr d}$ into cells indexed by $\Sigma_{m}\wr\Sigma_{d}$. In order to achive that, we first examine how far the group $\Sigma_{m}\wr\Sigma_{d}$ is from a Weyl group arising from a Tits system.
Definition 2.2.1.
A Tits system consists of the following quadruple $(G,B,N,W)$ in which $G$ is generated by its subgroups $B$ and $N$, $W:=N/(B\cap N)$ is well-defined and is generated by involutive elements $s_{i}(i\in I)$. Moreover, for all $s=s_{i}$, $w\in W$:
$sBw\subseteq BswB\cup BwB,\quad sB\neq Bs.$ (2.2.1)
Example 2.2.2.
Let $K$ be a field, $B_{n}:=B_{n}(K)$ be the standard Borel subgroup of $\mathrm{GL}_{n}:=\mathrm{GL}_{n}(K)$, and $N_{n}:=N_{n}(K)$ be the group of monomial matrices in $\mathrm{GL}_{n}$. Here we identify $K^{\times}\wr\Sigma_{n}\equiv N_{n}$, $(a_{1},\dots,a_{n},w)\mapsto\sum_{i=1}^{n}a_{i}E_{i,w^{-1}(i)}$.
  1. (i)
    The symmetric group $W=\Sigma_{n}$ is produced from the Tits system $(\mathrm{GL}_{n},B_{n},N_{n},\Sigma_{n})$.
  2. (ii)
    Let $\mathbb{Q}_{p}$ be the field of $p$-adic numbers, and $\mathbb{Z}_{p}$ be the ring of $p$-adic integers. The (non-extended) affine Weyl group $W=\Sigma^{\mathrm{aff}}_{n}$ is produced from the Tits system $(\mathrm{GL}_{n}(\mathbb{Q}_{p}),B,N_{n}(\mathbb{Q}_{p}),\Sigma^{\mathrm{aff}}_{n})$ where $B=\{(a_{ij})\in\mathrm{GL}_{n}(\mathbb{Z}_{p})~{}|~{}a_{ij}\in p\mathbb{Z}_{p}~{}\textup{if }i>j\}$.
In order to find a Tits-type system $(G,B,N,W)$ so that $W=\Sigma_{m}\wr\Sigma_{d}$, it makes sense to assume that
$\begin{split}N=N_{m\wr d}&:=\{A\equiv(a_{1},\dots,a_{md},w)\in N_{md}~{}|~{}w\in\Sigma_{m}\wr\Sigma_{d}\}\\ &=K^{\times}\wr(\Sigma_{m}\wr\Sigma_{d})=(K^{\times}\wr\Sigma_{m})\wr\Sigma_{d}=N_{m}\wr\Sigma_{d},\end{split}$ (2.2.2)
where $N_{m\wr d}$ is the group of monomial matrices corresponding to elements in $\Sigma_{m}\wr\Sigma_{d}$, regarded as a subgroup of $\Sigma_{md}$. Next, having $\Sigma_{m}\wr\Sigma_{d}\cong N_{m\wr d}/(N_{m\wr d}\cap B)$ in mind, we need to find a subgroup $B\subseteq\mathrm{GL}_{md}$ whose intersection with $N_{m\wr d}$ is equal to the subgroup of diagonal matrices in $\mathrm{GL}_{md}$. Hence, it makes sense to assume further
$B=B_{m}^{d}:=\{\textup{diag}(b_{1},\dots,b_{d})\in\mathrm{GL}_{md}~{}|~{}b_{i}\in B_{m}~{}\textup{for all}~{}i\}=B_{m}\wr\{1_{\Sigma_{d}}\},$ (2.2.3)
where $\textup{diag}(b_{1},\dots,b_{d})$ is the block matrix obtained by putting $b_{i}$’s in the diagonal. Thus, we are in a position to define our group $G$.
Definition 2.2.3.
Recall from (2.2.2)–(2.2.3) the subgroups $B_{m}^{d},N_{m\wr d}$ of $\mathrm{GL}_{md}$. Set
$G_{m\wr d}:=\langle B_{m}^{d},N_{m\wr d}\rangle\subseteq\mathrm{GL}_{md}.$
Now, we look back to our group $W=\Sigma_{m}\wr\Sigma_{d}$. While its generators $s_{1},\dots,s_{m-1},t_{1},\dots,t_{d-1}$ are indeed of order two and that $s_{i}$’s satisfy (2.2.1), the remaining generators $t_{j}$’s do not satisfy (2.2.1). In particular, $t_{j}B=Bt_{j}$ for all $j$.
In the following section, we will see that one can still obtain a Bruhat decomposition for $G_{m\wr d}$ via Iwahori’s generalized Tits system.
2.3. Iwahori’s Generalized Tits system
Iwahori showed that we can relax the conditions on a Tits system to obtain a large index set $W$ for the cells such that $W$ is a semidirect product of a Coxeter group $W_{0}$ and a group $\Omega$ consisting of certain elements $t$ such that $tB=Bt$.
Definition 2.3.1.
([I65]) Iwahori’s generalized Tits system is a quadruple $(G,B,N,W)$ in which $G$ is generated by its subgroups $B$ and $N$, $W:=N/(B\cap N)$ is well-defined. Moreover,
$W\cong W_{0}\rtimes\Omega,$ (2.3.1)
where $W_{0}$ is generated by involutive elements $s_{i}(i\in I)$, any element $t\in\Omega$ normalizes both $B$ and $\{s_{i}\}_{i\in I}$, and $Bt\neq B$ unless $t=1$. Finally, (2.2.1) holds for all $s=s_{i}$, $w\in W$.
The generalized Tits systems enjoy the following properties.
Lemma 2.3.2 ([I65]).
Assume that $(G,B,N,W)$ is a generalized Tits system. Then:
  1. (a)
    There is a Bruhat decomposition $G=\bigsqcup_{w\in W}BwB$.
  2. (b)
    $G_{0}:=BW_{0}B$ is a normal subgroup of $G$. Moreover, $(G_{0},B,N\cap G_{0},W_{0})$ is a Tits system. In particular, $W_{0}$ is a Coxeter group.
Then we have the following result.
Proposition 2.3.3.
The quadruple $(G_{m\wr d},B_{m}^{d},N_{m\wr d},\Sigma_{m}\wr\Sigma_{d})$ forms a generalized Tits system.
Proof.
Let $T_{m}:=B_{m}\cap N_{m}$. We first check that $T:=B_{m}^{d}\cap N_{m\wr d}=T_{m}^{d}$ is a normal subgroup of $N_{m\wr d}$. Recall the shorthand we adapt in (2.1.2). Indeed, for any $(\tau_{i})_{i}\in T$ and $\gamma=(g_{i})_{i}w\in N_{m\wr d}$,
$\gamma(\tau_{i})_{i}\gamma^{-1}=(g_{i})_{i}w(\tau_{i})_{i}w^{-1}(g_{i}^{-1})_{i}=(g_{i}\tau_{w^{-1}(i)})_{i}ww^{-1}(g_{i}^{-1})_{i}=(g_{i}\tau_{w^{-1}(i)}g^{-1}_{i})_{i}\in T,$ (2.3.2)
since each $g_{i}\in N_{m}$. Thus, $W=N_{m\wr d}/T\equiv(N_{m}\wr\Sigma_{d})/T_{m}^{d}=(N_{m}/T_{m})\wr\Sigma_{d}\cong\Sigma_{m}^{d}\rtimes\Sigma_{d}$.
Following the notations in Example 2.1.1, the semidirect product $W=W_{0}\rtimes\Omega$ can be described via
$W_{0}=N_{m}^{d}/T\cong\langle s_{i}\in\Sigma_{m}\wr\Sigma_{d}~{}|~{}i\in I\rangle,\quad\Omega=(\{1_{\Sigma_{m}}\}\wr\Sigma_{d})/T\cong\{t_{k}\in\Sigma_{m}\wr\Sigma_{d}~{}|~{}1\leq k<d\},$ (2.3.3)
where $I=[1,md]\setminus m\mathbb{Z}$, and $s_{i}=s_{a}^{(b)}$ (cf. (2.1.4)) if $i=(b-1)m+a$. It then follows from (2.1.6) that $\Omega$ normalizes $\{s_{i}\}_{i\in I}$.
In order to verify whether $t_{k}\in\Omega$ normalizes $B$, we write $t_{k}=\dot{t}_{k}T$ for a fixed representative $\dot{t}_{k}\in N_{m\wr d}$. Then, any element in $Bt_{k}$ is of the form $(b_{i})_{i}\dot{t}_{k}T_{m}^{d}$ for some $(b_{i})_{i}\in B_{m}^{d}$, and thus
$(b_{i})_{i}\dot{t}_{k}T_{m}^{d}=\dot{t}_{k}(b_{t_{k}(i)})_{i}T_{m}^{d}\in\dot{t}_{k}BT=\dot{t}_{k}TB=t_{k}B.$ (2.3.4)
That is, $tB=Bt$ for all $t\in\Omega$, by symmetry. Moreover, for $t\in\Omega$, $Bt\neq B$ unless $t=1$.
Finally, we verify that (2.2.1) holds for all $s=\sigma^{(j)}(\sigma\in\Sigma_{m})$ and $w=(w_{i})_{i}t\in W(t\in\Omega)$: Note that $sBw=\sigma^{(j)}B_{m}^{d}(w_{i})_{i}t=(\pi_{i}B_{m}w_{i})_{i}t$ where $\pi_{i}$ is identity except for that $\pi_{j}=\sigma$. Since $(\mathrm{GL}_{m},B_{m},N_{m},\Sigma_{m})$ is a Tits system, $\sigma B_{m}w_{j}\subseteq B_{m}\sigma w_{j}B_{m}\cup B_{m}w_{j}B_{m}$, and hence
$sBw\subseteq B\sigma^{(j)}(w_{i})_{i}Bt\cup B(w_{i})_{i}Bt=B\sigma^{(j)}(w_{i})_{i}tB\cup B(w_{i})_{i}tB=BswB\cup BwB.$ (2.3.5)
Also, $sB\neq Bs$ follows from that $\sigma B_{m}\neq B_{m}\sigma$. ∎
Corollary 2.3.4.
There is a Bruhat decomposition
$G_{m\wr d}=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}B^{d}_{m}wB^{d}_{m}.$ (2.3.6)
Proof.
It follows by combining Lemma (2.3.2)(a) and Proposition 2.3.3. ∎
Remark 2.3.5.
The exact argument in the proof of Proposition 2.3.3 can be used to prove that if $(G,B,N,W)$ is a Tits system, then $(\langle B^{d},N\wr\Sigma_{d}\rangle,B^{d},N\wr\Sigma_{d},W\wr\Sigma_{d})$ is a generalized Tits system. However, we did not find it relevant to consider these systems other than the case when $W=\Sigma_{m}$.
2.4. Lie Group Structure
Recall from Definition 2.2.3 that $G_{m\wr d}$ is a subgroup of the Lie group $\mathrm{GL}_{md}$. We now identify the (abstract) wreath product $\mathrm{GL}_{m}\wr\Sigma_{d}$ with the group $G_{m\wr d}$. The proposition below follows from a routine calculation:
Proposition 2.4.1.
The assignment below gives a group isomorphism $\mathrm{GL}_{m}\wr\Sigma_{d}\to G_{m\wr d}$:
$(g_{i})_{i}w\mapsto\mathrm{diag}(g_{1},\dots,g_{d})\Theta_{w}=\Theta_{w}\mathrm{diag}(g_{w(1)},\dots,g_{w(d)}),$ (2.4.1)
where $\mathrm{diag}(g_{1},\dots,g_{d})$ is the corresponding block diagonal matrix in $\mathrm{GL}_{md}$, and $\Theta_{w}\in\mathrm{GL}_{md}$ is the permutation matrix corresponding to $(1^{d},w)\in\Sigma_{m}\wr\Sigma_{d}\subseteq\Sigma_{md}$.
For example, when $d=3$, the tuple $(A_{1},A_{2},A_{3},(1~{}2~{}3)=|2~{}3~{}1|)$ is sent to
$\begin{pmatrix}A_{1}&0&0\\ 0&A_{2}&0\\ 0&0&A_{3}\end{pmatrix}\begin{pmatrix}0&0&I_{m}\\ I_{m}&0&0\\ 0&I_{m}&0\end{pmatrix}=\begin{pmatrix}0&0&A_{1}\\ A_{2}&0&0\\ 0&A_{3}&0\end{pmatrix}=\begin{pmatrix}0&0&I_{m}\\ I_{m}&0&0\\ 0&I_{m}&0\end{pmatrix}\begin{pmatrix}A_{2}&0&0\\ 0&A_{3}&0\\ 0&0&A_{1}\end{pmatrix}.$ (2.4.2)
Remark 2.4.2.
The subgroup $G_{m\wr d}\subseteq\mathrm{GL}_{md}$ is a matrix Lie group in the sense that, for any sequence $\{A_{i}\}_{i\geq 1}$ in $G_{m\wr d}$ that converges to some matrix $A$, either $A\in G_{m\wr d}$ or $A$ is not invertible. Thus, $G_{m\wr d}$ is a closed subgroup of $\mathrm{GL}_{md}$. Applying the closed subgroup theorem, one gets that
$\textup{Lie}(G_{m\wr d})=\{X\in\mathfrak{gl}_{md}~{}|~{}e^{tX}\in G_{m\wr d}\textup{~{}for all~{}}t\in\mathbb{R}\}=\mathfrak{gl}_{m}^{d}.$ (2.4.3)
We remark that the corresponding Lie algebra is not the Lie subalgebra $\mathfrak{gl}_{m}\wr\Sigma_{d}\subseteq\mathfrak{gl}_{md}$.
3. Flag Varieties
3.1. Complete Flag Varieties
For now, let $\mathrm{GL}_{n}=\mathrm{GL}_{n}(K)$, $B_{n}=B_{n}(K)$ with corresponding Lie algebras $\mathfrak{gl}_{n}=\textup{Lie}(\mathrm{GL}_{n})$ and $\mathfrak{b}_{n}=\textup{Lie}(B_{n})$, respectively. It is well-known (see [CG97, §3]) that the following maps are bijections:
$\mathrm{GL}_{n}/B_{n}\to\mathfrak{B}_{n}\to\mathcal{F}l_{n},\quad gB_{n}\mapsto g\mathfrak{b}_{n}g^{-1}\mapsto gF^{\textup{std}},$ (3.1.1)
where $\mathfrak{B}_{n}:=\{\textup{solvable Lie subalgebra }\mathfrak{a}\subseteq\mathfrak{gl}_{n}~{}|~{}\dim\mathfrak{a}={\dim\mathfrak{b}_{n}}\}\hookrightarrow\mathbb{P}(\bigwedge^{{\dim\mathfrak{b}_{n}}}\mathfrak{gl}_{n})$ is the projective variety consisting of all Borel subalgebras of $\mathfrak{gl}_{n}$, and $\mathcal{F}l_{n}\hookrightarrow\prod_{i}\mathbb{P}^{*}(\bigwedge^{i}K^{n})$ is the (projective) complete flag variety, in which the standard flag in $K^{n}$ is denoted by $F^{\textup{std}}_{\bullet}$.
In our case, it turns out that one can only establish a bijection $G_{m\wr d}/B^{d}_{m}\cong\mathcal{F}l_{m\wr d}$ for some projective variety $\mathcal{F}l_{m\wr d}$ to be constructed. Note that, as we point out in Subsection 1.3, the variety $\mathfrak{B}_{m\wr d}$ of all Borel subalgebras in the wreath setting is still $\mathfrak{B}^{d}_{m}$, which does not in bijection with $G_{m\wr d}\cong\mathcal{F}l_{m\wr d}$. For example, when $m=d=2$, we have
$(g_{1},g_{2},s_{1})(\mathfrak{b}_{2},\mathfrak{b}_{2},id)(g_{1},g_{2},s_{1})^{-1}=(g_{1}\mathfrak{b}_{2}g^{-1}_{1},g_{2}\mathfrak{b}_{2}g^{-1}_{2},id)=(g_{1},g_{2},id)(\mathfrak{b}_{2},\mathfrak{b}_{2},id)(g_{1},g_{2},id)^{-1}$
where $(g_{1},g_{2},s_{1}),(g_{1},g_{2},id)\in\mathrm{GL}_{2}\wr\Sigma_{2}$ and $(\mathfrak{b}_{2},\mathfrak{b}_{2},id)\in\mathfrak{B}^{2}_{2}$.
3.2. Wreath Flags
We first construct a subvariety $\mathcal{F}l_{m\wr d}\subseteq\mathcal{F}l_{md}$ which affords a $G_{m\wr d}$-equivariant bijection with the homogeneous space $G_{m\wr d}/B^{d}_{m}$. We will show that $\mathcal{F}l_{m\wr d}$ can be identified with the wreath product $\mathcal{F}l_{m}\wr\Sigma_{d}$.
Fix a basis $\{e_{1},\dots,e_{md}\}$ of $K^{md}$, and let $V_{i}=\textup{Span}_{K}\{e_{(i-1)m+1},...,e_{im}\}$ for $1\leq i\leq d$.
Definition 3.2.1.
Denote the set which we call the wreath flag variety by
$\mathcal{F}l_{m\wr d}:=\{F_{\bullet}\in\mathcal{F}l_{md}~{}|~{}F_{im}=V_{w(1)}\oplus\dots\oplus V_{w(i)}\textup{ for all }1\leq i\leq d,\textup{ for some }w\in\Sigma_{d}\}.$ (3.2.1)
For any complete flag $F_{\bullet}$ in $V_{1}$ (i.e., $F_{\bullet}\in\mathcal{F}l_{m}$), denote by ${}^{i}F_{\bullet}$ the corresponding complete flag in $V_{i}$ obtained by shifting the indices of basis elements by $(i-1)m$. In particular, the corresponding standard flag in each $V_{i}$ is
${}^{i}F^{\textup{std}}_{\bullet}=(0\subseteq Ke_{(i-1)m+1}\subseteq Ke_{(i-1)m+1}\oplus Ke_{(i-1)m+2}\subseteq\dots\subseteq V_{i}).$ (3.2.2)
Proposition 3.2.2.
  1. (a)
    $\mathcal{F}l_{m}\wr\Sigma_{d}\equiv\mathcal{F}l_{m\wr d}$ as sets via $(F^{1}_{\bullet},\dots,F^{d}_{\bullet},w)\mapsto\mathcal{F}_{\bullet}=\mathcal{F}_{\bullet}(F^{1}_{\bullet},\dots,F^{d}_{\bullet},w)$, where
    $\mathcal{F}_{im+j}=V_{w(1)}\oplus\dots\oplus V_{w(i)}\oplus({}^{w(i+1)}F^{w(i+1)}_{j}),\quad(0\leq i<d,1\leq j<m).$ (3.2.3)
  2. (b)
    There is a natural identification $G_{m\wr d}/B^{d}_{m}\to\mathcal{F}l_{m\wr d}$, $(g_{i})_{i}wB^{d}_{m}\mapsto(g_{1}F^{std}_{\bullet},...,g_{d}F^{std}_{\bullet},w)$.
  3. (c)
    Under the identifications in Part (a) and in Proposition 2.4.1, the $G_{m\wr d}$-action on $\mathcal{F}l_{m\wr d}$ given by
    $(g_{i})_{i}w(F^{1}_{\bullet},...,F^{d}_{\bullet},\sigma)=(g_{1}F^{w^{-1}(1)}_{\bullet},...,g_{d}F^{w^{-1}(d)}_{\bullet},w\sigma)$ (3.2.4)
    is compatible with the $\mathrm{GL}_{md}$-action on $\mathcal{F}l_{md}$.
Proof.
Part (a) follows from a direct verification. Part (b) follows from Part (a). For Part (c), it suffices to show that
$\Theta_{w}\mathcal{F}_{\bullet}(F^{1}_{\bullet},\dots,F^{d}_{\bullet},\sigma)$ $=\mathcal{F}_{\bullet}(F^{w^{-1}(1)}_{\bullet},\dots,F^{w^{-1}(d)}_{\bullet},w\sigma),$ (3.2.5)
$\mathrm{diag}(g_{1},\dots,g_{d})\mathcal{F}_{\bullet}(F_{\bullet}^{1},...,F_{\bullet}^{d},\sigma)$ $=\mathcal{F}_{\bullet}(g_{1}F_{\bullet}^{1},...,g_{d}F_{\bullet}^{d},\sigma).$ (3.2.6)
Write $(F^{1}_{\bullet},...,F^{d}_{\bullet},\sigma)=(h_{i})_{i}\sigma(F^{\textup{std}}_{\bullet},...,F^{\textup{std}}_{\bullet},id)$ for some $(h_{i})_{i}\sigma\in G_{m\wr d}$. Then, a standard calculation shows that $\mathrm{diag}(h_{1},\dots,h_{d})\Theta_{\sigma}\mathcal{F}_{\bullet}(F_{\bullet}^{std},...,F_{\bullet}^{std},id)=\mathcal{F}_{\bullet}(F^{1}_{\bullet},...,F^{d}_{\bullet},\sigma)$. Thus, (3.2.5) is given by
$\begin{split}\Theta_{w}\mathcal{F}_{\bullet}(F^{1}_{\bullet},\dots,F^{d}_{\bullet},\sigma)&=\Theta_{w}\mathrm{diag}(h_{1},\dots,h_{d})\Theta_{\sigma}\mathcal{F}_{\bullet}(F_{\bullet}^{std},...,F_{\bullet}^{std},id)\\ &=\mathrm{diag}(h_{w^{-1}(1)},\dots,h_{w^{-1}(d)})\Theta_{w}\Theta_{\sigma}\mathcal{F}_{\bullet}(F_{\bullet}^{std},...,F_{\bullet}^{std},id)\\ &=\mathrm{diag}(h_{w^{-1}(1)},\dots,h_{w^{-1}(d)})\Theta_{w\sigma}\mathcal{F}_{\bullet}(F_{\bullet}^{std},...,F_{\bullet}^{std},id)\\ &=\mathcal{F}_{\bullet}(F^{w^{-1}(1)}_{\bullet},\dots,F^{w^{-1}(d)}_{\bullet},w\sigma).\end{split}$ (3.2.7)
(3.2.6) follows from a similar verification. ∎
3.3. Geometric Bruhat decomposition
Next, we give a geometric proof of the Bruhat decomposition of the wreath flag variety $\mathcal{F}l_{m\wr d}$ based on the Bialynicki-Birula decomposition.
Let $X$ be a smooth complex projective variety with an algebraic $\mathbb{C}^{*}$-action. Assume that the set $X^{\mathbb{C}^{*}}$ of $\mathbb{C}^{*}$-fixed points of $X$ is finite. For $w\in X^{\mathbb{C}^{*}}$, denote by $X_{w}=\{x\in X\ |\ \lim_{z\rightarrow 0}z\cdot x=w\}$ the attracting set. Since $\mathbb{C}^{*}$ fixes $w$, there is a natural $\mathbb{C}^{*}$-action on the tangent space $T_{w}X$. Set $T^{+}_{w}X:=\bigoplus_{n\in\mathbb{Z}_{>0}}T_{w}X(n)$, where $T_{w}X(n)\coloneqq\{x\in T_{w}X\ |\ z\cdot x=z^{n}x\ \forall z\in\mathbb{C}^{*}\}$.
Proposition 3.3.1 (Bialynicki-Birula decomposition).
Let $W=X^{\mathbb{C}^{*}}$ be the (finite) set of $\mathbb{C}^{*}$-fixed points of a smooth complex projective variety $X$. Then,
  1. (a)
    The attracting sets form a decomposition $X=\bigsqcup_{w\in W}X_{w}$ into smooth locally closed subvarieties;
  2. (b)
    Each attractung set $X_{w}$ is isomorphic to $T_{w}X_{w}=T^{+}_{w}X$ as algebraic varieties. The isomophism commutes with the $\mathbb{C}^{*}$-action.
The first step is to choose a suitable $\mathbb{C}^{*}$-action on our wreath flag variety $X=\mathcal{F}l_{m\wr d}\equiv G_{m\wr d}/B^{d}_{m}$. Fix a maximal torus $T^{d}_{m}\subseteq B^{d}_{m}\subseteq\mathrm{GL}_{m}^{d}$, and let it act on $X$ by
$(t_{1},...,t_{d})\cdot(g_{1},...,g_{d},w)B^{d}_{m}=(t_{1}g_{1},...,t_{d}g_{d},w)B^{d}_{m}.$ (3.3.1)
The next step is to analyze the set of $T^{d}_{m}$-fixed points.
Lemma 3.3.2.
There is a bijection $\mathcal{F}l_{m\wr d}^{T_{m}^{d}}\to\Sigma_{m}\wr\Sigma_{d}$.
Proof.
Let $(g_{i})_{i}wB^{d}_{m}$ be a $T^{d}_{m}$-fixed point in $\mathcal{F}l_{m\wr d}\equiv G_{m\wr d}/B^{d}_{m}$. Then, for any $(t_{i})_{i}\in T^{d}_{m}$, $(t_{i})_{i}(g_{i})_{i}wB^{d}_{m}=(g_{i})_{i}wB^{d}_{m}$, and hence
$B^{d}_{m}=((g_{i})_{i}w)^{-1}(t_{i})_{i}(g_{i})_{i}wB^{d}_{m}=(g^{-1}_{w(i)})_{i}w^{-1}(t_{i}g_{i})_{i}wB^{d}_{m}=(g^{-1}_{w(i)}t_{w(i)}g_{w(i)})_{i}B^{d}_{m}.$ (3.3.2)
Thus, for each $1\leq i\leq d$, we have $g^{-1}_{i}t_{i}g_{i}B_{m}=B_{m}$, or equivalently, $T_{m}\subseteq g_{i}B_{m}g^{-1}_{i}$. Therefore, there is a bijection
$\mathcal{F}l_{m\wr d}^{T_{m}^{d}}\to\{\textup{Borel subgroups of }\mathrm{GL}_{m}^{d}\textup{ containing }T_{m}^{d}\}\times\Sigma_{d},\quad(g_{i})_{i}wB^{d}_{m}\mapsto\textstyle{\left(\prod_{i=1}^{m}g_{i}B_{m}g^{-1}_{i},w\right)}.$ (3.3.3)
Since $g_{i}B_{m}g_{i}^{-1}\subseteq\mathrm{GL}_{m}$ is a Borel subgroup which contains $T_{m}$, the set of $T^{d}_{m}$-fixed points of $\mathcal{F}l_{m\wr d}$ is the same as the set of Borel subgroups of $\mathrm{GL}_{m}^{d}$ containing $T^{d}_{m}$ times the extra copy $\Sigma_{d}$. The lemma concludes from the fact that there is a bijection between the set of all Borel subgroups of $\mathrm{GL}_{m}$ containing $T_{m}$ and the symmetric group $\Sigma_{m}$.
Theorem 3.3.3 (Geometric Bruhat decomposition).
Let $X=\mathcal{F}l_{m\wr d}\equiv G_{m\wr d}/B^{d}_{m}$ be the wreath flag variety. Then,
  1. (a)
    There is a $\mathbb{C}^{*}$-action on $X$ such that the fixed-point set $X^{\mathbb{C}^{*}}$ is in bijection with $\Sigma_{m}\wr\Sigma_{d}$.
  2. (b)
    $G_{m\wr d}/B^{d}_{m}=\bigsqcup_{\sigma\in\Sigma_{m}\wr\Sigma_{d}}X_{\sigma}$, where $X_{\sigma}$ is precisely the cell $B_{m}^{d}\sigma B^{d}_{m}/B_{m}^{d}$.
Proof.
[…]
Combining (a) and Proposition 3.3.1, we obtain that $X$ decomposes into locally closed subvarieties of the form $X_{\sigma}\cong T^{+}_{\sigma}\mathcal{F}l_{m\wr d}$ ($\sigma\in\Sigma_{m}\wr\Sigma_{d}$). Write $\sigma=(g_{i})_{i}w$ for some $g_{i}\in\Sigma_{m},w\in\Sigma_{d}$. Then,
$T_{\sigma}X\cong T_{(g_{1},...,g_{d})}\mathrm{GL}^{d}_{m}/B^{d}_{m}\times\{w\}={\textstyle\prod_{i=1}^{d}}T_{g_{i}}(\mathrm{GL}_{m}/B_{m})\times\{w\}.$ (3.3.4)
By the proof of [CG97, Corollary 3.1.12], $T^{+}_{g_{i}}(\mathrm{GL}_{m}/B_{m})\cong B_{m}g_{i}B_{m}/B_{m}$ for all $i$. Thus,
$T^{+}_{\sigma}\mathcal{F}l_{m\wr d}\cong{\textstyle\prod_{i=1}^{d}}T^{+}_{g_{i}}(\mathrm{GL}_{m}/B_{m})\times\{w\}\cong{\textstyle\prod_{i=1}^{d}}(B_{m}g_{i}B_{m}/B_{m})\times\{w\}\cong B^{d}_{m}\sigma B^{d}_{m}/B^{d}_{m}.$ (3.3.5)
4. A Steinberg variety
In this section, we introduce a Steinberg variety $Z_{m\wr d}$, in order to prepare a realization of the group algebra of $\Sigma_{m}\wr\Sigma_{d}$ via the top Borel-Moore homology of $Z_{m\wr d}$. Thus, we analyze the irreducible components of $Z_{m\wr d}$ in terms of the conormal bundles. We remark that the argument presented here is not a consequence of [CG97, Proposition 3.3.4] since the obvious wreath vairants of Borel subalgebras are not in bijection with the wreath flags.
4.1. Nilpotent Orbits
As observed in Remark 2.4.2, we should consider nilpotent cones for the Lie algebra $\mathfrak{gl}_{m}^{d}$, instead of a wreath version. Let $\mathcal{N}_{m}$ be the variety of all nilpotent elements in $\mathfrak{gl}_{m}$. The $\mathrm{GL}_{m}$-orbits in $\mathcal{N}_{m}$ are parametrized by the set $\Pi_{m}$ of partitions $\nu$ of $m\in\mathbb{Z}_{\geq 0}$ (write $\nu\vdash m$ for short).
Let $G_{m\wr d}$ act on $\mathcal{N}^{d}_{m}$ by
$\gamma\cdot(n_{i})_{i}:=\gamma(n_{i})_{i}\gamma^{-1}=(g_{i}n_{w^{-1}(i)}g^{-1}_{i})_{i},$ (4.1.1)
for $\gamma=(g_{i})_{i}w\in G_{m\wr d}$ and $(n_{i})_{i}\in\mathcal{N}_{m}^{d}$. Note that (4.1.1) is compatible with the $\mathrm{GL}_{md}$-action on $\mathcal{N}_{md}$ via the embedding $\mathcal{N}_{m}^{d}\hookrightarrow\mathcal{N}_{md},(n_{i})_{i}\mapsto\mathrm{diag}(n_{1},\dots,n_{d})$. Let $\Sigma_{d}$ act on $\Pi_{m}^{d}$ from the right by place permutations. Thus, the $G_{m\wr d}$-orbits in $\mathcal{N}_{m}^{d}$ are parametrized by the $\Sigma_{d}$-orbits in the multipartitions $(\lambda_{1},\dots,\lambda_{d})$ where $\lambda_{i}\vdash m$ for all $i$, i.e.,
$G_{m\wr d}\backslash\mathcal{N}_{m}^{d}\overset{1:1}{\longleftrightarrow}\{(\lambda_{1},...,\lambda_{d})\cdot\Sigma_{d}~{}|~{}\lambda_{i}\vdash m\textup{ for all }i\}.$ (4.1.2)
Remark 4.1.1.
Under the identification $\mathcal{F}l_{m\wr d}\equiv\mathcal{F}l_{m}\wr\Sigma_{d}$, any $\mathcal{F}\in\mathcal{F}l_{m\wr d}\subseteq\mathcal{F}l_{md}$ is identified with some element of the form $(F^{i}_{\bullet})_{i}w\in\mathcal{F}l_{m}\wr\Sigma_{d}$. It is standard to check that $\mathcal{F}$ is nilpotent in $\mathcal{F}l_{md}$ if and only if $n_{i}F^{i}_{j}\subseteq F^{i}_{j-1}$ for all $i,j$.
4.2. A Variant of Springer Resolution
Denote the Springer resolution of type A by $\mu_{m}:\widetilde{\mathcal{N}}_{m}\to\mathcal{N}_{m}$, $(u,F_{\bullet})\mapsto u$, where
$\widetilde{\mathcal{N}}_{m}=T^{*}\mathcal{F}l_{m}\equiv\{(u,F_{\bullet})\in\mathcal{N}_{m}\times\mathcal{F}l_{m}\ |\ u(F_{i})\subseteq F_{i-1}\textup{ for all }i\}.$ (4.2.1)
Instead of the usual Springer resolution $\mu_{m}^{d}:\widetilde{\mathcal{N}}_{m}^{d}\to\mathcal{N}_{m}^{d}$, we consider the following Springer resolution:
$\mu_{m\wr d}:\widetilde{\mathcal{N}}_{m\wr d}\rightarrow\mathcal{N}_{m}^{d},\ (\nu,\mathcal{F}_{\bullet})\mapsto\nu,\quad\textup{where}$ (4.2.2)
$\widetilde{\mathcal{N}}_{m\wr d}:=T^{*}\mathcal{F}l_{m\wr d}\equiv\{((n_{i})_{i},(F^{i}_{\bullet})_{i}w))\in\mathcal{N}_{m}^{d}\times\mathcal{F}l_{m\wr d}\ |\ n_{i}F^{i}_{j}\subseteq F^{i}_{j-1}\ \text{for \ all}\ i,\ j\}.$ (4.2.3)
Let $\mathfrak{n}_{m}^{d}\subseteq\mathfrak{b}_{m}^{d}$ be the nilpotent radical. Recall $G_{m\wr d}$ acts on $\mathcal{N}_{m}^{d}$ and $\mathcal{F}l_{m\wr d}$ by (4.1.1) and (3.2.4), respectively, and hence $G_{m\wr d}$ acts on $\widetilde{\mathcal{N}}_{m\wr d}:=T^{*}\mathcal{F}l_{m\wr d}$. Then, we have the following analog of [CG97, Corollary 3.1.33] regarding $G_{m\wr d}$-equivariant vector bundles.
Proposition 4.2.1.
As $G_{m\wr d}$-equivariant vector bundles, $G_{m\wr d}\times_{B_{m}^{d}}\mathfrak{n}_{m}^{d}\cong T^{*}\mathcal{F}l_{m\wr d}$ via
$(\gamma,\nu)\mapsto(\gamma\cdot\nu,\ \gamma(F^{std}_{\bullet},...,F^{std}_{\bullet},1_{\Sigma_{d}})),\quad\gamma\in G_{m\wr d},\nu\in\mathfrak{n}_{m}^{d}.$ (4.2.4)
Proof.
For injectivity, suppose that $(\gamma_{1}\cdot\nu_{1},\ \gamma_{1}(F^{std}_{\bullet},...,F^{std}_{\bullet},1_{\Sigma_{d}}))=(\gamma_{2}\cdot\nu_{2},\ \gamma_{2}(F^{std}_{\bullet},...,F^{std}_{\bullet},1_{\Sigma_{d}}))$ for some $(\gamma_{i},\nu_{i})\in G_{m\wr d}\times_{B_{m}^{d}}\mathfrak{n}_{m}^{d}$. Then, $\gamma_{1}^{-1}\gamma_{2}\in\textup{Stab}_{G_{m\wr d}}((F^{std}_{\bullet},...,F^{std}_{\bullet},1_{\Sigma_{d}}))=B^{d}_{m}$, and thus $\gamma_{2}=\gamma_{1}b$ for some $b\in B^{d}_{m}$. Moreover,
$\gamma_{2}\cdot\nu_{2}=\gamma_{2}\nu_{2}\gamma_{2}^{-1}=\gamma_{1}b\nu_{2}b^{-1}\gamma_{1}^{-1}=\gamma_{1}\nu_{1}\gamma_{1}^{-1}=\gamma_{1}\cdot\nu_{1},$ (4.2.5)
which implies that $b\nu_{2}b^{-1}=\nu_{1}$. Therefore, $(\gamma_{1},\nu_{1})=(\gamma_{2}b^{-1},b\nu_{2}b^{-1})=(\gamma_{2},\nu_{2})$.
For surjectivity, assume that $((n_{i})_{i},(F^{i}_{\bullet})_{i}w))\in T^{*}\mathcal{F}l_{m\wr d}$. Then, there exists $\gamma=(g_{i})_{i}w\in G_{m\wr d}$ such that $\gamma(F^{std}_{\bullet},...,F^{std}_{\bullet},1_{\Sigma_{d}})=(F^{i}_{\bullet})_{i}w))$. The condition $n_{i}F^{i}_{j}\subseteq F^{i}_{j-1}$ implies that $g_{i}^{-1}n_{i}g_{i}\in\mathfrak{n}_{m}$ for all $i$. Thus, $((g_{i})_{i}w,(g^{-1}_{w(i)}n_{w(i)}g_{w(i)})_{i})$ lies in the inverse image, and we are done. ∎
We identify $\mathfrak{gl}_{m}^{d}\cong\mathfrak{gl}_{m}^{d,*}$ by an invariant bilinear form on $\mathfrak{gl}_{m}^{d}$, and let $(\mathfrak{b}^{d}_{m})^{\perp}\subseteq\mathfrak{gl}_{m}^{d,*}$ be the image of $\mathfrak{n}^{d}_{m}$ under the identification. Then, we obtain the following analog of [CG97, Proposition 1.4.9].
Corollary 4.2.2.
There is a $G_{m\wr d}$-equivariant isomorphism $T^{*}\mathcal{F}l_{m\wr d}\cong G_{m\wr d}\times_{B_{m}^{d}}(\mathfrak{b}^{d}_{m})^{\perp}$.
4.3. Wreath Steinberg Varieties
Using our unconventional Springer resolution $\widetilde{\mathcal{N}}_{m\wr d}:=T^{*}\mathcal{F}l_{m\wr d}$, we introduce the following Steinberg varieties whose irreducible components are conormal bundles indexed by $\Sigma_{m}\wr\Sigma_{d}$.
Definition 4.3.1.
Define the wreath Steinberg variety by $Z_{m\wr d}:=\widetilde{\mathcal{N}}_{m\wr d}\times_{\mathcal{N}_{m}^{d}}\widetilde{\mathcal{N}}_{m\wr d}$, i.e,
$Z_{m\wr d}=\{((\nu,\mathcal{F}_{\bullet}),(\nu^{\prime},\mathcal{F}^{\prime}_{\bullet}))\in\widetilde{\mathcal{N}}_{m\wr d}\times\widetilde{\mathcal{N}}_{m\wr d}\ |\ \nu=\nu^{\prime}\}.$ (4.3.1)
It then follows from the Bruhat decomposition (2.3.6) that
$(G_{m\wr d}/B^{d}_{m})\times(G_{m\wr d}/B^{d}_{m})=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}\mathcal{O}(w),\quad\mathcal{O}(w):=G_{m\wr d}\cdot((F^{\textup{std}}_{\bullet})_{i},w(F^{\textup{std}}_{\bullet})_{i}).$ (4.3.2)
Thus, for $w\in\Sigma_{m}\wr\Sigma_{d}$, there is a short exact sequence via the identification (Proposition 3.2.2(b)):
$0\rightarrow Y_{w}\rightarrow T^{*}(\mathcal{F}l_{m\wr d}\times\mathcal{F}l_{m\wr d})\rightarrow T^{*}\mathcal{O}(w)\rightarrow 0,$ (4.3.3)
where $Y_{w}:=T^{*}_{\mathcal{O}(w)}(\mathcal{F}l_{m\wr d}\times\mathcal{F}l_{m\wr d})$ is the cornormal bundle. On the other hand, there is a natural projection $\pi:Z_{m\wr d}\rightarrow G_{m\wr d}/B^{d}_{m}\times G_{m\wr d}/B^{d}_{m}$ given by
$Z_{m\wr d}\hookrightarrow T^{*}\mathcal{F}l_{m\wr d}\times T^{*}\mathcal{F}l_{m\wr d}\xrightarrow{\pi^{\prime}\times\pi^{\prime}}(G_{m\wr d}/B^{d}_{m})\times(G_{m\wr d}/B^{d}_{m}),$ (4.3.4)
where $\pi^{\prime}:T^{*}\mathcal{F}l_{m\wr d}\to G_{m\wr d}/B_{m}^{d}$ is the evident projection. In the following, we show that $Z_{w}:=\pi^{-1}(\mathcal{O}(w))$ gives an alternative description for the conormal bundle $Y_{w}$.
Lemma 4.3.2.
For $w\in\Sigma_{m}\wr\Sigma_{d}$, let $Z_{w}\coloneqq\pi^{-1}(\mathcal{O}(w))$. Then, $Z_{w}=Y_{w}$.
Proof.
An arbitrary geometric point in the orbit $\mathcal{O}(w)$ is of the form $s=(F^{(1)},F^{(2)})$ where
$F^{(j)}=(g^{(j)}_{i})_{i}\sigma_{1}(F^{std}_{\bullet},...,F^{std}_{\bullet}),\quad(j=1,2,\ g^{(j)}_{i}\in\mathrm{GL}_{m},\sigma_{j}\in\Sigma_{d})$ (4.3.5)
such that $((g^{(1)}_{i})_{i}\sigma_{1})^{-1}(g^{(2)}_{i})_{i}\sigma_{2}\in B^{d}_{m}wB^{d}_{m}$. Then, by Corollary 4.2.2, the fiber $Y_{w,s}$ of the cornormal bundle of $\mathcal{O}(w)$ at the point $s$ consists of elements of the form $((n^{(1)},n^{(2)}),(F^{(1)},F^{(2)}))$, where $n^{(1)},n^{(2)}\in\mathfrak{gl}_{m}^{d,*}$ and $(n^{(1)},n^{(2)})$ annihilates the tangent space $T_{s}\mathcal{O}(w)$. Since $\textup{Stab}_{G_{m\wr d}}(F^{std}_{\bullet},...,F^{std}_{\bullet})=B^{d}_{m}$, the stabilizer of $F^{(j)}$ in $G_{m\wr d}$ is a Borel subgroup $B_{j}$ of $G_{m\wr d}$, i.e.,
$\textup{Stab}_{G_{m\wr d}}(F^{(j)})=(g^{(j)}_{i})_{i}\sigma_{j}B^{d}_{m}((g^{(j)}_{i})_{i}\sigma_{j})^{-1}=B_{j}.$ (4.3.6)
Therefore, the orbit $\mathcal{O}(w):=G_{m\wr d}\cdot s$ is given by
$\mathcal{O}(w)=G_{m\wr d}/(\textup{Stab}(F^{(1)})\cap\textup{Stab}(F^{(2)}))=G_{m\wr d}/(B_{1}\cap B_{2}).$ (4.3.7)
Write $\mathfrak{b}_{j}=\mathop{\mathrm{Lie}}B_{j}$. Since $G_{m\wr d}/(B_{1}\cap B_{2})=\{g(B_{1}\cap B_{2})\ |\ g\in G_{m\wr d}\}\cong\{(gB_{1},gB_{2})\ |\ g\in G_{m\wr d}\}$,
$T_{s}\mathcal{O}(w)\cong T_{s}G_{m\wr d}/(B_{1}\cap B_{2})\cong T_{s}\{(gB_{1},gB_{2})\ |\ g\in G_{m\wr d}\}=\{(x\mathfrak{b}_{1},x\mathfrak{b}_{2})\ |\ x\in\mathfrak{gl}^{d}_{m}\}.$ (4.3.8)
Now, the condition that $(n^{(1)},n^{(2)})$ annihilates the tangent space $T_{s}\mathcal{O}(w)$ is equivalent to $\langle n^{(1)},x\rangle+\langle n^{(2)},x\rangle=0$ for all $x\in\mathfrak{gl}^{d}_{m}$, which implies that $n^{(1)}=-n^{(2)}$. Thus, we have
$Y_{w,s}=\{((n^{(1)},n^{(2)}),(F^{(1)},F^{(2)}))\in T_{s}^{*}(\mathcal{F}l_{m\wr d}\times\mathcal{F}l_{m\wr d})\ |\ n^{(1)}=-n^{(2)}\}.$ (4.3.9)
Finally, $Y_{w,s}$ coincides with the fiber $Z_{w,s}$ thanks to the sign isomorphism $T^{*}(\mathcal{F}l_{m\wr d}\times\mathcal{F}l_{m\wr d})\xrightarrow{sign}T^{*}\mathcal{F}l_{m\wr d}\times T^{*}\mathcal{F}l_{m\wr d}=\widetilde{\mathcal{N}}_{m\wr d}\times\widetilde{\mathcal{N}}_{m\wr d}$, $((n^{(1)},n^{(2)}),(F^{(1)},F^{(2)}))T\mapsto((n^{(1)},F^{(1)}),(-n^{(2)},F^{(2)}))$. ∎
Proposition 4.3.3.
There is a decomposition $Z_{m\wr d}=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}Y_{w}$. Moreover, each irreducible component of $Z_{m\wr d}$ is a closure of $Y_{w}$ for a unique $w\in\Sigma_{m}\wr\Sigma_{d}$.
Proof.
Observe that $Z_{m\wr d}\subseteq\mathcal{N}_{m}^{d}\times G_{m\wr d}/B^{d}_{m}\times G_{m\wr d}/B^{d}_{m}=\mathcal{N}_{m}^{d}\times\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}\mathcal{O}(w)$. Thus
$Z_{m\wr d}=Z_{m\wr d}\cap(\mathcal{N}_{m}^{d}\times\textstyle{\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}\mathcal{O}(w)})=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}Z_{m\wr d}\cap(\mathcal{N}_{m}^{d}\times\mathcal{O}(w))=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}Z_{w}.$
The proposition then follows from Lemma 4.3.2. ∎
5. Borel-Moore Homology of Wreath Steinberg Varieties
5.1. Borel-Moore Homology
An introduction to theory of Borel-Moore homology can be found in [CG97, §2.6–7]. In particular, we will use the following result:
Proposition 5.1.1.
Let $M$ be a smooth complex manifold, $\pi:M\to N$ be a proper map onto a variety, and let $Z=M\times_{N}M:=\{(m,m^{\prime})\in M^{2}~{}|~{}\pi(m)=\pi(m^{\prime})\}$. Denote by $H^{\mathrm{BM}}(Z;K)$ the group of Borel-Moore homology with coefficients in a field $K$ of characteristic 0, and by $H^{\mathrm{BM}}_{\mathrm{top}}(Z;K)$ its top degree homologies. Then,
  1. (1)
    [CG97, Corollary 2.7.41, 2.7.48] The group $H^{\mathrm{BM}}_{\mathrm{top}}(Z;K)$ has a structure of an associative unital $K$-algebra.
  2. (2)
    [CG97, Lemma 2.7.49] Assume that the set of irreducible components $\{L_{w}\}_{w\in W}$ of $Z$ is indexed by a finite set $W$. If all the components have the same dimension, then the fundamental classes $[L_{w}]$, for all $w\in W$, form a basis of the algebra $H^{\mathrm{BM}}_{\mathrm{top}}(Z;K)$.
Following [CG97, §3], the above-mentioned results apply to the the Steinberg variety $Z_{m}$ for $\mathrm{GL}_{m}$. Write $H^{\mathrm{BM}}_{\mathrm{top}}(Z):=H^{\mathrm{BM}}_{\mathrm{top}}(Z;\mathbb{Q})$. It is further proved therein that the top Borel-Moore homology with rational coefficients gives a geometric realization of the group algebra of $\Sigma_{m}$ as below:
Proposition 5.1.2.
  1. (1)
    [CG97, Theorem 3.4.1, Claim 3.4.13] There is an algebra isomorphism $\mathbb{Q}[\Sigma_{m}]\to H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m}),w\mapsto[\Lambda^{0}_{w}]$, where $[\Lambda^{0}_{w}]$ is a certain fundamental class obtained by taking specialization.
  2. (2)
    [CG97, Lemma 3.4.14] For $w\in\Sigma_{m}$, let $Y^{\prime}_{w}$ be the conormal bundle to the orbit $\mathcal{O}(w)\in\mathrm{GL}_{m}\backslash(\mathcal{F}l_{m}\times\mathcal{F}l_{m})$. Then, with respect to the Bruhat order on $\Sigma_{m}$,
    $[\Lambda^{0}_{w}]\in[Y^{\prime}_{w}]+{\textstyle\sum_{y<w}}\mathbb{Q}[Y^{\prime}_{y}].$ (5.1.1)
Remark 5.1.3.
Following [CG97, §3.4], for each $w\in\Sigma_{m}$ and a fixed regular vector $h$ in the abstract Cartan subalgebra $\mathfrak{H}$ of $\mathfrak{gl}_{m}$, there is a fiber $\Lambda_{w}^{h}:=(\nu\times\nu)^{-1}(w.h,h)$, where $\nu$ is the projection $\widetilde{\mathfrak{gl}}_{m}\to\mathfrak{H}$. That is, $\Lambda_{w}^{h}=\{(x,\mathfrak{b},\mathfrak{b}^{\prime})\in\mathfrak{gl}_{m}\times\mathfrak{B}_{m}\times\mathfrak{B}_{m}~{}|~{}x\in\mathfrak{b}\cap\mathfrak{b}^{\prime},\ \nu(x,\mathfrak{b})=w.h,\ \nu(x,\mathfrak{b}^{\prime})=h\}$. By taking the specialization map (see [CG97, (2.6.30)]), one defines a class $[\Lambda_{w}^{0}]$ that is not the class of some subvariety, and does not depend on the choice of $h$.
It is tempting to mimic this approach to construct the desirable classes for the generators $t_{k}$’s in $\Sigma_{m}\wr\Sigma_{d}$ (see Example 2.1.1). However, such a specialization map produces classes $[\Lambda_{t_{k}}^{0}]$ that do not sit inside $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$, in general. We have to construct explicitly classes in $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$ that play the same role as the elements $t_{k}$’s in the group algebra.
5.2. A Geometric Realization for $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]$
It follows from Definition 4.3.1 that $Z^{d}_{m}\subseteq Z_{m\wr d}$, and hence by Proposition 5.1.2(1),
$H^{\mathrm{BM}}_{\mathrm{top}}(Z^{d}_{m})=H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m})^{\otimes d}\cong\mathbb{Q}[\Sigma_{m}]^{\otimes d}\subseteq H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d}).$ (5.2.1)
By a slight abuse of notation, for any class $[C]\in H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m})$, we denote by $[C]^{(j)}$ the corresponding class in $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$ via (2.1.4) and (5.2.1). Recall also the conormal bundle $[Y_{t_{k}}]$ defined in (4.3.3). We are now in a position to prove the analog of the geometric realization theorem for the group algebra of $\Sigma_{m}\wr\Sigma_{d}$ in terms of the wreath flag variety $Z_{m\wr d}$.
Theorem 5.2.1.
The assignment $s_{i}^{(j)}\mapsto[\Lambda_{s_{i}}^{0}]^{(j)},\ t_{k}\mapsto[Y_{t_{k}}]$ defines an algebra isomorphism $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]\cong H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$.
Proof.
By Proposition 5.1.1, the algebras $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]$ and $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$ have the same dimension. It suffices to check the following relations hold, for $21\leq j\leq d-1,|j-i|\geq 2,1\leq k\leq d-$ 2:
(quadratic relations) $[Y_{t_{j}}]^{2}=[\Lambda^{0}_{1}],$ (5.2.2)
(wreath relations) $[Y_{t_{j}}]*[\Lambda^{0}_{(w_{1},\dots,w_{d})}]=[\Lambda^{0}_{(w_{1},\dots,w_{d}).s_{j}}]*[Y_{t_{j}}]\quad((w_{1},\dots,w_{d})\in\Sigma_{m}^{d}),$ (5.2.3)
(braid relations) $[Y_{t_{k}}]*[Y_{t_{k+1}}]*[Y_{t_{k}}]=[Y_{t_{k+1}}]*[Y_{t_{k}}]*[Y_{t_{k+1}}],$ (5.2.4)
$[Y_{t_{i}}]*[Y_{t_{j}}]=[Y_{t_{j}}]*[Y_{t_{i}}].$ (5.2.5)
To verify (5.2.2) and (5.2.3), it suffices to consider the “rank one” case. That is, consider $t=t_{1}\in\Sigma_{m}\wr\Sigma_{2}$, we need to show that
$[Y_{t}]\ast[Y_{t}]=[\Lambda^{0}_{1}].$ (5.2.6)
Denote by $\pi_{ij}$ the natural projections from $\widetilde{\mathcal{N}}_{m\wr 2}\times_{\mathcal{N}_{m}^{2}}\widetilde{\mathcal{N}}_{m\wr 2}\times_{\mathcal{N}_{m}^{2}}\widetilde{\mathcal{N}}_{m\wr 2}$ onto the $(i,j)$th components. Then,
$\pi^{-1}_{12}(Y_{t})\cap\pi_{23}^{-1}(Y_{t})=\{(\nu,\mathcal{F},\mathcal{F}^{\prime},\mathcal{F}^{\prime\prime})\in\mathcal{N}_{m}^{2}\times\mathcal{F}l_{m\wr 2}^{3}\ |\ (\mathcal{F},\mathcal{F}^{\prime}),(\mathcal{F}^{\prime},\mathcal{F}^{\prime\prime})\in\mathcal{O}(t)\}.$ (5.2.7)
Applying $\pi_{13}$ to (5.2.7), we obtains the following set-theoretic composition:
$Y_{t}\circ Y_{t}=\pi_{13}(\pi^{-1}_{12}(Y_{t})\cap\pi_{23}^{-1}(Y_{t}))=\{(\nu,\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{N}_{m}^{2}\times\mathcal{F}l_{m\wr 2}^{2}\ |\ (\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{O}(1)\},$ (5.2.8)
which is equal to $Y_{1}=\Lambda^{0}_{1}$. Since $\pi_{13}:\pi^{-1}_{12}(Y_{t})\cap\pi_{23}^{-1}(Y_{t})\rightarrow Y_{t}\circ Y_{t}$ is bijective, its fiber $F$ is a point. That is, the Euler characteristic $\chi(F)=1$. By [CG97, Theorem 2.7.26(b)], $[Y_{t}]\ast[Y_{t}]=\chi(F)[Y_{t}\circ Y_{t}]=[\Lambda^{0}_{e}]$, which proves (5.2.6).
For (5.2.3), it suffices prove the following, for any $s=s_{i}\in\Sigma_{m}$:
$[Y_{t}]\ast[\Lambda^{0}_{s^{(1)}}]=[\Lambda^{0}_{s^{(2)}}]\ast[Y_{t}].$ (5.2.9)
By Proposition 5.1.2(2), $[\Lambda^{0}_{s}]=[Y^{\prime}_{s}]+q$ for some $q\in\mathbb{Q}$, and hence for $j=1$ or 2, $[\Lambda^{0}_{s^{(j)}}]=[Y_{s^{(j)}}]+q$ for the same $q$. Therefore, (5.2.9) is equivalent to $[Y_{t}]\ast[Y_{s^{(1)}}]=[Y_{s^{(2)}}]\ast[Y_{t}]$. Note that
$Y_{t}\circ Y_{s^{(1)}}=\pi_{13}(\pi^{-1}_{12}(Y_{t})\cap\pi_{23}^{-1}(Y_{s^{(1)}}))=\{(\nu,\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{N}_{m}^{2}\times\mathcal{F}l_{m\wr 2}^{2}\ |\ (\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{O}(ts^{(1)})\}.$ (5.2.10)
Since the map $\pi_{13}:\pi^{-1}_{12}(Y_{t})\cap\pi_{23}^{-1}(Y_{s^{(1)}})\rightarrow Y_{t}\circ Y_{s^{(1)}}$ is bijective, its fiber $F$ is a point, and hence
$[Y_{t}]\ast[Y_{s^{(1)}}]=[Y_{t}\circ Y_{s^{(1)}}]=[Y_{ts^{(1)}}]=[Y_{s^{(2)}t}].$ (5.2.11)
A similar calculation then shows that $Y_{t}\circ Y_{s^{(1)}}=Y_{s^{(2)}}\circ Y_{t}$. Thus, we prove (5.2.9).
It remains to prove the braid relations (5.2.4) – (5.2.5) for the “rank three case” in $\Sigma_{m}\wr\Sigma_{4}$, i.e.,
$[Y_{t_{1}}]\ast[Y_{t_{2}}]\ast[Y_{t_{1}}]=[Y_{t_{2}}]\ast[Y_{t_{1}}]\ast[Y_{t_{2}}],\quad{[Y_{t_{1}}]\ast[Y_{t_{3}}]=[Y_{t_{3}}]\ast[Y_{t_{1}}]}.$ (5.2.12)
Note that
$Y_{t_{1}}\circ Y_{t_{2}}=\{(\nu,\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{N}_{m}^{4}\times\mathcal{F}l_{m\wr 4}^{2}\ |\ (\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{O}(t_{1}t_{2})\}=Y_{t_{1}t_{2}}.$ (5.2.13)
It is standard to check that $\pi_{13}:\pi^{-1}_{12}(Y_{t_{1}})\cap\pi_{23}^{-1}(Y_{t_{2}})\rightarrow Y_{t_{1}}\circ Y_{t_{2}}$ is bijective. Thus, $[Y_{t_{1}}]\ast[Y_{t_{2}}]=[Y_{t_{1}}\circ Y_{t_{2}}]=[Y_{t_{1}t_{2}}]$. Consequently,
$Y_{t_{1}t_{2}}\circ Y_{t_{1}}=\{(\nu,\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{N}_{m}^{4}\times\mathcal{F}l_{m\wr 4}^{2}\ |\ (\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{O}(t_{1}t_{2}t_{1})\}=Y_{t_{1}t_{2}t_{1}}.$ (5.2.14)
Again, the map $\pi_{13}:\pi^{-1}_{12}(Y_{t_{1}t_{2}})\cap\pi_{23}^{-1}(Y_{t_{1}})\rightarrow Y_{t_{1}t_{2}}\circ Y_{t_{2}}$ is bijective, and hence $([Y_{t_{1}}]\ast[Y_{t_{2}}])\ast[Y_{t_{1}}]=[Y_{t_{1}t_{2}}]\ast[Y_{t_{1}}]=[Y_{t_{1}t_{2}}\circ Y_{t_{1}}]=[Y_{t_{1}t_{2}t_{1}}]$.
Since $Y_{t_{1}}\circ Y_{t_{3}}=\{(\nu,\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{N}_{m}^{4}\times\mathcal{F}l_{m\wr 4}^{2}\ |\ (\mathcal{F},\mathcal{F}^{\prime\prime})\in\mathcal{O}(t_{1}t_{3})\}=Y_{t_{1}t_{3}}$, a similar argument shows that $[Y_{t_{1}}]\ast[Y_{t_{3}}]=[Y_{t_{1}}\circ Y_{t_{3}}]=[Y_{t_{1}t_{3}}]$. Then, (5.2.12) follows by symmetry. The proof is complete. ∎
6. Irreducible representations
In this section, we obtain a geometric classification of the simples of $\Sigma_{m}\wr\Sigma_{d}$ as a corollary of Theorem 5.2.1. We further identify the classification with the well-known classification via Clifford theory.
6.1. Ginzburg’s Lagrangian Construction
We aim to apply the classification theorem [CG97, Theorem 3.5.7] in Chriss-Ginzburg. First, we recall the required assumptions listed in [CG97, §3.5]:
Definition 6.1.1.
Let $G$ be an algebraic group with a Borel subgroup $B$. We call a morphism $\mu:\widetilde{\mathcal{N}}\rightarrow\mathcal{N}$ of $G$-variety a Springer resolution if the following conditions hold:
  1. (1)
    $\widetilde{\mathcal{N}}$ is smooth.
  2. (2)
    $\mathcal{N}$ has finitely many $G$-orbits.
  3. (3)
    $\mu:\widetilde{\mathcal{N}}\rightarrow\mathcal{N}$ is $G$-equivariant and proper.
  4. (4)
    (dimension property) For each $x\in\mathcal{N}$, all irreducible components of $\mathfrak{B}_{x}:=\mu^{-1}(x)$ have the same dimension given by $\dim\mathfrak{B}_{x}=\dim(G/B)-\frac{1}{2}\dim(G\cdot x)$.
Assume for now this section that $\mu$ is a Springer resolution. Let $Z:=\widetilde{\mathcal{N}}{\times}_{\mathcal{N}}\widetilde{\mathcal{N}}$ be the Steinberg variety. It follows from [CG97, 2.7.40] that each $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})$ has a left and a right $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$-module structure, denoted by $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}$ and $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{R}$, respectively.
For any finite-dimensional left module $V$ of $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$, denote by $V^{\vee}$ the right $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$-module with underlying space $V^{*}:=\mathrm{Hom}_{\mathbb{Q}}(V,\mathbb{Q})$ on which the action is given by
$(f\cdot a):v\mapsto f(a\cdot v),\quad a\in H^{\mathrm{BM}}_{\mathrm{top}}(Z),\ f\in V^{\vee},\ v\in V.$ (6.1.1)
For $x\in\mathcal{N}$, let $G(x)$ be the centralizer of $x$ in $G$, $G^{0}(x)$ be the identity component, and $C(x):=G/G^{0}(x)$ be the component group. Then, ${H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})$ is a module over $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$. Denote the set of all irreducible complex representations of $C(x)$ that occurs in $\mathbb{C}\otimes_{\mathbb{Q}}{H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})$, up to isomorphism, by
$C(x)^{\wedge}:=\{[L]~{}|~{}L\in{\mathop{\mathrm{Irr}}}\mathbb{C}[C(x)],[\mathbb{C}\otimes_{\mathbb{Q}}H^{\mathrm{BM}}(\mathfrak{B}_{x}):L]\neq 0\}.$ (6.1.2)
For each $\psi\in C(x)^{\wedge}$, let ${H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})_{\psi}$ be the $\psi$-isotypic component of ${H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})$.
Proposition 6.1.2 ([CG97, Theorem 3.5.7]).
Assume that $\mu$ is a Springer resolution in the sense of Definition 6.1.1, and that $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$ is semisimple. If, for all $x\in\mathcal{N}$, the isomorphism ${H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})_{R}\cong{H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})_{L}^{\vee}$ of right $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$-modules is compatible with their respective $C(x)$-actions, then the complete set of isomorphism classes of irreducible $H^{\mathrm{BM}}_{\mathrm{top}}(Z)$-modules over $\mathbb{C}$ is given by $\{{H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})_{\psi}~{}|~{}[x,\psi]\in\mathbb{I}\}$, where
$\mathbb{I}:=\{G\textup{-conjugacy class }[x,\psi]\textup{ of }(x,\psi)~{}|~{}x\in\mathcal{N},\psi\in C(x)^{\wedge}\}.$ (6.1.3)
We remark that, by [CG97, Claim 3.5.5], the left $C(x)$-module structure of ${H^{\mathrm{BM}}_{\mathrm{top}}}(\mathfrak{B}_{x})_{L}^{\vee}$ is given by
${(g\cdot{\check{v}})(v)}={\check{v}}(g^{-1}\cdot v),\quad\textup{where}\ g\in C(x),\ {\check{v}}\in H^{\mathrm{BM}}(\mathfrak{B}_{x})_{L}^{\vee},\ v\in H^{\mathrm{BM}}(\mathfrak{B}_{x})_{L}.$ (6.1.4)
6.2. A Springer-type Classification of Irreducibles
Let $G=G_{m\wr d}$, $B=B^{d}_{m}$, and recall the wreath Springer resolution $\mu=\mu_{m\wr d}:\widetilde{\mathcal{N}}_{m\wr d}\rightarrow\mathcal{N}_{m}^{d}$ from (4.2.2). For $x=(n_{i})_{i}\in\mathcal{N}^{d}_{m}$, the set theoretic description for Springer fiber is given by
$\mathfrak{B}_{x}=\{(F^{i}_{\bullet})_{i}w\in\mathcal{F}l_{m\wr d}\ |\ n_{i}F^{i}_{j}\subseteq F^{i}_{j-1}\ \text{for all}\ i,j\}.$ (6.2.1)
Proposition 6.2.1.
The wreath Springer resolution $\mu_{m\wr d}:T^{*}\mathcal{F}l_{m\wr d}\rightarrow\mathcal{N}_{m}^{d}$ is a Springer resolution in the sense of Definition 6.1.1.
Proof.
By Propositions 2.4.1 and 3.2.2 (a), there are identifications $G_{m\wr d}\equiv\mathrm{GL}_{m}\wr\Sigma_{d}$ and $\mathcal{F}l_{m\wr d}=\mathcal{F}l_{m}\wr\Sigma_{d}$, respectively. Thus, $G_{m\wr d}$ and $T^{*}\mathcal{F}l_{m\wr d}$ are disjoint unions of $d!$-copies of $\mathrm{GL}^{d}_{m}$ and $T^{*}\mathcal{F}l_{m}^{d}$, respectively, and hence (1) holds. Next, (2) follows from the orbit decomposition (4.1.2).
For (3), the $G_{m\wr d}$-equivariance follows from the fact that
$\mu_{m\wr d}(\gamma\cdot(x,\mathcal{F}))=\mu_{m\wr d}(\gamma\cdot x,\gamma\cdot\mathcal{F})=\gamma\cdot x=\gamma\cdot\mu_{m\wr d}(x,\mathcal{F}),$ (6.2.2)
for all $\gamma\in G_{m\wr d}$, $(x,\mathcal{F})\in\widetilde{\mathcal{N}}_{m\wr d}$. For properness, it follows from the fact that $\mu_{m\wr d}$ is the restriction of the projection $\mathfrak{gl}_{md}\times\mathcal{F}l_{md}\rightarrow\mathfrak{gl}_{md}$.
For the dimension property (4), let $x=(n_{i})_{i}\in\mathcal{N}_{m}^{d}$. Note that the fiber $\mathfrak{B}_{x}$ is a disjoint union of $d!$ copies of $\prod_{i=1}\mathfrak{B}_{n_{i}}$. Next, the orbit $\mathbb{O}:=G_{m\wr d}\cdot x$ has dimension equals to the sum of dimensions of all $\mathrm{GL}_{m}\cdot n_{i}$. Finally, $\dim(G_{m\wr d}/B_{m}^{d})=d\dim(\mathrm{GL}_{m}/B_{m})$, and hence
$\dim\mathfrak{B}_{x}={\textstyle\sum_{i=1}^{d}}\dim\mathfrak{B}_{n_{i}}=d\dim(\mathrm{GL}_{m}/B_{m})-{\textstyle\frac{1}{2}\sum_{i=1}^{d}}(\mathrm{GL}_{m}\cdot n_{i})=\dim(G_{m\wr d}/B_{m}^{d})-{\textstyle\frac{1}{2}}\dim\mathbb{O}.$ (6.2.3)
Note that switching factors in $Z_{m\wr d}$ defines an involution, which induces an algebra anti-involution $T$ on $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$ via $c\mapsto c^{T}$. Consequently, as left $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$-modules, $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}=H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{R}^{T}$, where the latter is the module with underlying space $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{R}$ on which $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$ acts by
$c\cdot v\coloneq v\cdot c^{T},\quad\textup{where}\ v\in H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x}),\ c\in H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d}).$ (6.2.4)
In order to show that the assumptions in Proposition 6.1.2 hold for $\mu_{m\wr d}$, we need the following lemma regarding $T$.
Lemma 6.2.2.
Under the isomorphism $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})\cong\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$, the anti-involution $T$ corresponds to the anti-involution on $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$ via $w\mapsto w^{-1}$ for all $w\in\Sigma_{m}\wr\Sigma_{d}$.
Proof.
It suffices to consider the generators $[\Lambda^{0}_{s}]$ and $[\Lambda^{0}_{t}]$, for some $s=s_{i}^{(j)}$ and some $t=t_{k}\in\Sigma_{m}\wr\Sigma_{d}$. It follows from [CG97, Lemma 3.6.11] that $(\Lambda^{0}_{s})^{T}=\Lambda^{0}_{s^{-1}}$. Moreover, since $\Lambda^{0}_{t}=Y_{t}$ and $t=t^{-1}$,
$\begin{split}(Y_{t})^{T}&=\{(x,\mathcal{F},(F^{i}_{\bullet})_{i}w)\ |\ (\mathcal{F},(F^{i}_{\bullet})_{i}w)=xw((F^{std}_{\bullet},...,F^{std}_{\bullet},t),(F^{std}_{\bullet},...,F^{std}_{\bullet},id))\}\\ &=\{(x,\mathcal{F},(F^{i}_{\bullet})_{i}w)\ |\ (\mathcal{F},(F^{i}_{\bullet})_{i}w)=xwt((F^{std}_{\bullet},...,F^{std}_{\bullet},id),(F^{std}_{\bullet},...,F^{std}_{\bullet},t))\}\\ &=Y_{t}=Y_{t^{-1}}.\end{split}$ (6.2.5)
We are now in a position to prove the classification theorem for simple modules over $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$.
Theorem 6.2.3.
The complete set of isomorphism classes of simple $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-modules is given by $\{H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{\psi}~{}|~{}[x,\psi]\in I^{G}_{m\wr d}\}$, where $I^{G}_{m\wr d}:=\{G_{m\wr d}\textup{-conjugacy class }[x,\psi]~{}|~{}x\in\mathcal{N}^{d}_{m},\ \psi\in C(x)^{\wedge})\}$.
Proof.
Note that $H^{\mathrm{BM}}_{\mathrm{top}}(Z_{m\wr d})$ is semisimple by Theorem 5.2.1. Thanks to Propositions 6.1.2 and 6.2.1, it suffices to show that there is an isomorphism $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}\cong(H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}^{\vee})^{T}$ that is compatible with the respective $C(x)$-actions, for all $x\in\mathcal{N}_{m}^{d}$.
Since $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}$ is finite dimensional, by Lemma 6.2.2, there is a isomorphism $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}\to{(H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}^{\vee})^{T}},v\mapsto v^{*}$ of left $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-modules Therefore, for each $g\in C(x),v\in H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{L}$, $(g\cdot v)^{*}(u)=v^{*}(g^{-1}\cdot u)=(g\cdot v^{*})(u)$ for all $u\in H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})$. Thus, all the assumptions are verified and we are done.
6.3. Clifford Theory
We take a detour to review the well-known classification of irreducible modules over the group algebra $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$ using Clifford theory, following [CST14, §2.6.1] (see also [CR81]).
Recall that $\Pi_{m}$ is the set of partitions $\nu$ of $m\in\mathbb{Z}_{\geq 0}$. Let $\Pi=\bigcup_{m\geq 0}\Pi_{m}$. Let $|\nu|\in\mathbb{Z}_{\geq 0}$ be the number partitioned by $\nu\in\Pi$. For $\nu\vdash m$, let $S^{\nu}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{m}]$ be the corresponding (irreducible) Specht module.
For any map $\lambda:\Pi_{m}\to\Pi$ (equivalently, a multi-partition $\lambda$ of ${}^{\#}\Pi_{m}$ components), define a map $|\lambda|:\Pi_{m}\to\mathbb{Z}_{\geq 0}$ by setting $|\lambda|(\nu):=|\lambda(\nu)|$ for all $\nu\vdash m$. Each map $\gamma:\Pi_{m}\to\mathbb{Z}_{\geq 0}$ is associated with a Young subgroup $\Sigma_{\gamma}:=\prod_{\nu\vdash m}\Sigma_{\gamma(\nu)}\subseteq\Sigma_{d}$. Set
$I^{C}_{m\wr d}:=\{\lambda:\Pi_{m}\to\Pi~{}|~{}\textstyle{\sum_{\nu\vdash m}}|\lambda(\nu)|=d\},\quad J_{m\wr d}:=\{\gamma:\Pi_{m}\to\mathbb{Z}_{\geq 0}~{}|~{}\textstyle{\sum_{\nu\vdash m}}\gamma(\nu)=d\}.$ (6.3.1)
Pick $\lambda\in I^{C}_{m\wr d}$ and let $\gamma=|\lambda|$. We further define irreducible modules
$S^{\lambda}:=\bigotimes_{\nu\vdash m}S^{\lambda(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma}],\quad S^{\gamma}:=\bigotimes_{\nu\vdash m}(S^{\nu})^{\otimes\gamma(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{m}^{d}].$ (6.3.2)
Denote the extension of $S^{\gamma}$ and the inflation of $S^{\lambda}$ to their inertia group $\Sigma_{m}\wr\Sigma_{\gamma}$ by $\widetilde{S^{\gamma}}\in\mathop{\mathrm{Mod}}\mathbb{C}[\Sigma_{m}\wr\Sigma_{\gamma}]$ and $\mathop{\mathrm{Infl}}S^{\lambda}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{m}\wr\Sigma_{\gamma}]$, respectively. That is, $\widetilde{S^{\gamma}}=S^{\gamma}$ and $\mathop{\mathrm{Infl}}S^{\lambda}=S^{\lambda}$ as vector spaces, while the actions of additional group elements are given by, respectively, place permutations on tensor factors, and the trivial action.
Proposition 6.3.1.
The complete list of non-isomorphic irreducible representations of $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$ is given by
$\left\{{\mathrm{Ind}}_{\Sigma_{m}\wr\Sigma_{|\lambda|}}^{\Sigma_{m}\wr\Sigma_{d}}(\widetilde{S^{|\lambda|}}\otimes\mathop{\mathrm{Infl}}S^{\lambda})~{}\middle|~{}\lambda\in I^{C}_{m\wr d}\right\}.$
6.4. Characterization of the Index Set
In this subsection, we give an explicit characterization for the index set of the simple $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-modules from the geometric approach in Theorem 6.2.3 and provide an identification to the index set in the algebraic approach from the above subsection.
From Theorem 6.2.3, each irreducible is indexed by a $G_{m\wr d}$-conjugacy class $[x,\psi]$ for some $x=(n_{i})_{i}\in\mathcal{N}_{m}^{d}$ and $\psi\in C(x)^{\wedge}$. Each $n_{i}\in\mathcal{N}_{m}$ has Jordan type $x_{i}\vdash m$. Equivalently, $x$ is associated with a map $\gamma(x)\in J_{m\wr d}$ given by
$\gamma(x)(\nu):={}^{\#}\{i\in\{1,\dots,d\}~{}|~{}x_{i}=\nu\}.$ (6.4.1)
Proposition 6.4.1.
Let $x\in\mathcal{N}^{d}_{m}$, and let $\gamma=\gamma(x)\in J_{m\wr d}$ as in (6.4.1). Then, the component group $C(x)$ is isomorphic to the Young subgroup $\Sigma_{\gamma(x)}$.
Proof.
Fix an element $x=(n_{i})_{i}\in\mathcal{N}^{d}_{m}$. There are $\gamma(\nu)$ entries $n_{i}$’s in $x$ that are of Jordan type $\nu$, for each $\nu$. The centralizer of $x$ in $G_{m\wr d}$ is given by
$\begin{split}C_{G_{m\wr d}}(x)&=\{(g_{i})_{i}w\in G_{m\wr d}\ |\ (g_{i}n_{w^{-1}(i)}g^{-1}_{i})_{i}=(n_{i})_{i}\ w\in\Sigma_{d}\}\\ &=\{(g_{i})_{i}w\in G_{m\wr d}\ |\ (g_{i}n_{w^{-1}(i)}g^{-1}_{i})_{i}=(n_{i})_{i},\ w\in\Sigma_{\gamma}\},\end{split}$ (6.4.2)
since $n_{i}$ is conjugate to $n_{j}$ if and only if $x_{i}=x_{j}$. Moreover, the identity of $G_{m\wr d}$ lies in $C_{G_{m\wr d}}(x)$.
For now, we abbreviate the identity component $C_{G_{m\wr d}}(x)^{0}$ by $C^{0}$. We are going to show that the component group $C(x):=C_{G_{m\wr d}}(x)/C^{0}$ is isomorphic to $\Sigma_{\gamma}:=\prod_{\nu\vdash m}\Sigma_{\gamma(\nu)}$. First, we assume that $n_{1}$, $\dots$, $n_{\gamma(x_{1})}$ are of the same type $x_{1}\vdash m$, and then construct corresponding elements in $C(x)$ that generate a copy of $\Sigma_{\gamma(x_{1})}$.
To be precise, consider elements of the form $(g_{1},\cdots,g_{\gamma(x_{1})},1,\cdots,1)t_{k}\in C_{G_{m\wr d}}(x)$ for some $g_{i}\in\mathrm{GL}_{m}$ and $t_{k}\in\Sigma_{\gamma(x_{1})}$. For all $11\leq j\leq\gamma(x_{1})-$ 1, let $r_{j}\in\mathrm{GL}_{m}$ be such that $r_{j}n_{j+1}r_{j}^{-1}=n_{j}$. Thus, each equality $g_{i}n_{t_{k}^{-1}(i)}g^{-1}_{i}=n_{i}$ appearing in (6.4.2) becomes one of the following: $g_{k}n_{k+1}g^{-1}_{k}=r_{k}n_{k+1}r_{k}^{-1}$, $g_{k+1}n_{k}g^{-1}_{k+1}=r_{k}^{-1}n_{k}r_{k}$, or $g_{i}n_{i}g_{i}^{-1}=n_{i}$ if $i\neq k,k+1$. Equivalently, there exists $g^{\prime}_{i}\in C_{\mathrm{GL}_{m}}(n_{i})$ for all $1\leq i\leq\gamma(x_{1})$ such that
$g_{k}=r_{k}g^{\prime}_{k}\quad g_{k+1}=r_{k}^{-1}g^{\prime}_{k+1}\quad g_{i}=g^{\prime}_{i}\ \textup{if }i\neq k,k+1.$ (6.4.3)
Let $\sigma_{k}:=(g_{1},\cdots,g_{\gamma(x_{1})},1^{d-\gamma(x_{1})})t_{k}C^{0}\in C(x)$ for $11\leq k\leq\gamma(\nu)-$ 1. Then,
$\begin{split}\sigma_{k}&=(g^{\prime}_{1},\dots,g^{\prime}_{k-1},r_{k}g^{\prime}_{k+1},r_{k}^{-1}g^{\prime}_{k},g^{\prime}_{k+2},\dots,g^{\prime}_{\gamma(x_{1})},1^{d-\gamma(x_{1})})t_{k}C^{0}\\ &=(1^{k-1},r_{k},r_{k}^{-1},1^{d-k-1})t_{k}(g^{\prime}_{1},\dots,g^{\prime}_{\gamma(x_{1})},1^{d-\gamma(x_{1})})C^{0}=(1^{k-1},r_{k},r_{k}^{-1},1^{d-k-1})t_{k}C^{0},\end{split}$ (6.4.4)
and thus $\sigma_{k}$ is of order 2. Next, we verify the braid relations for $\sigma_{1},\dots,\sigma_{\gamma(x_{1})}$. We may assume that $d=4=\gamma(\nu_{1})$. For adjacent braids, consider $\sigma_{1}:=(r_{1},r^{-1}_{1},1,1)t_{1}C^{0}$, $\sigma_{2}:=(1,r_{2},r^{-1}_{2},1)t_{2}C^{0}$. The braid relation follows from the fact that
$\begin{split}\sigma_{1}\sigma_{2}\sigma_{1}&=(r_{1}r_{2},r^{-1}_{1},r^{-1}_{2},1)t_{1}t_{2}(r_{1},r^{-1}_{1},1,1)t_{1}C^{0}=(r_{1}r_{2},1,r_{2}^{-1}r^{-1}_{1},1)t_{1}t_{2}t_{1}C^{0},\\ \sigma_{2}\sigma_{1}\sigma_{2}&=(r_{1},r_{2},r_{2}^{-1}r_{1}^{-1},1)t_{2}t_{1}(1,r_{2},r^{-1}_{2},1)t_{2}C^{0}=(r_{1}r_{2},1,r_{2}^{-1}r_{1}^{-1},1)t_{2}t_{1}t_{2}C^{0}.\end{split}$ (6.4.5)
For non-adjacent braids (if any), consider $\sigma_{3}:=(1,1,r_{3},r_{3}^{-1})t_{3}C^{0}$, and then
$\sigma_{1}\sigma_{3}=(r_{1},r^{-1}_{1},r_{3},r_{3}^{-1})t_{1}t_{3}C^{0}=\sigma_{3}\sigma_{1}.$ (6.4.6)
Therefore, $C(x)\cong\Sigma_{\gamma(x_{1})}\times C_{G_{m\wr(d-\gamma(x_{1}))}}(n_{\gamma(x_{1})}+1,\dots n_{d})$. It then follows from an induction on $d$ that $C(x)\cong\Sigma_{\gamma}$, and we are done. ∎
Recall $C(x)^{\wedge}$ from (6.1.2). Next, we show that $C(x)^{\wedge}$ coincides with all irreducibles over $\mathbb{C}[\Sigma_{\gamma(x)}]$.
Proposition 6.4.2.
Let $x\in\mathcal{N}_{m}^{d}$. Then, $C(x)^{\wedge}=\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma(x)}]$.
Proof.
It suffices to show that $C(x)^{\wedge}\supseteq\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma(x)}]$. That is, given $x\in\mathcal{N}_{m}^{d}$ and $\lambda\in I^{C}_{m\wr d}$ such that $|\lambda|=\gamma(x)$,
$[\mathbb{C}\otimes_{\mathbb{Q}}H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x}):S^{\lambda}]\neq 0.$ (6.4.7)
Since $C(x)$ is a product of Weyl groups of type A, (6.4.7) holds if and only if $[H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x};\mathbb{C}):S^{\lambda}]\neq 0$.
It follows from (6.2.1) and the identification $\mathcal{F}l_{m\wr d}\equiv\mathcal{F}l_{m}\wr\Sigma_{d}$ in Proposition 3.2.2 (a) that
$H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x};\mathbb{C})\cong\bigoplus_{w\in\Sigma_{d}}w\biggl{(}\bigotimes_{\nu\vdash m}(S^{\nu})^{\otimes\gamma(x)(\nu)}\biggr{)}=\bigoplus_{w\in\Sigma_{d}}wS^{\gamma(x)},$ (6.4.8)
where $\Sigma_{d}$ acts by tensor factor permutations (since $\Sigma_{\nu\vdash m}\gamma(x)(\nu)=d$), and thus gives the explicit $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-module structure on $H^{BM}_{top}(\mathfrak{B}_{x},\mathbb{C})$.
Let $\lambda\in I^{C}_{m\wr d}$ be a multi-partition such that $|\lambda|=\gamma(x)$. Recall from (6.3.2) the Specht module $S^{\lambda}$ over $\mathbb{C}[\Sigma_{\gamma(x)}]$. Via the inclusion $\Sigma_{\gamma(x)}\subseteq\Sigma_{d}\subseteq\Sigma_{m}\wr\Sigma_{d}$, $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x};\mathbb{C})$ affords a $\mathbb{C}[\Sigma_{\gamma(x)}]$-module structure. Moreover, from (6.4.8), $H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x};\mathbb{C})$ contains the subspace $\bigoplus_{w\in\Sigma_{\gamma(x)}}wS^{\gamma(x)}$, which is a regular representation of $\Sigma_{\gamma(x)}$. Since the regular representation contains all irreducible representations, we are done.
As a consequence, we can now identify the irreducibles arising from Clifford theory with those arising from Springer theory.
Corollary 6.4.3.
There assignment $\lambda\mapsto[x(\lambda),S^{\lambda}]$ defines a bijection $\Psi:I^{C}_{m\wr d}\to I^{G}_{m\wr d}$, where $x(\lambda)$ is a fixed element in $\mathcal{N}_{m}^{d}$ such that $\gamma(x(\lambda))=|\lambda|$.
Conversely, for any $[x,\psi]\in I^{G}_{m\wr d}$, $\psi\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma(x)}]$ is of the form $S^{\lambda}:=\bigotimes_{\nu\vdash m}S^{\lambda(\nu)}$ for some $\lambda\in I^{C}_{m\wr d}$. Then, the inverse map is given by $[x,\psi]\mapsto\lambda$.
6.5. Special Cases
First, we consider an extreme case that $m=2$ so that the wreath product $\Sigma_{2}\wr\Sigma_{d}$ is a Weyl group of type $B_{d}$.
Example 6.5.1.
The index set $I^{C}_{2\wr d}=\{\lambda:\{(2),(1,1)\}\to\Pi~{}|~{}|\lambda((2))|+|\lambda((1,1))|=d\}$ is in bijection with the set of all bipartitions of $d$ via $\lambda\mapsto(\lambda((2)),\lambda((1,1)))$.
For the geometric index set $I^{G}_{2\wr d}$, note that any $x\in\mathcal{N}_{2}^{d}$ has Jordan type $(2)^{a}(1,1)^{d-a}$ for some $a$. The component group $C_{G_{2\wr d}}(x)$ is then isomorphic to the Young subgroup $\Sigma_{a}\times\Sigma_{d-a}$, and hence any $\psi\in C(x)^{\wedge}$ is of the form $S^{\lambda^{\prime}}\otimes S^{\lambda^{\prime\prime}}$ for some $\lambda^{\prime}\vdash a,\lambda^{\prime\prime}\vdash d-a$. Therefore, $\Psi$ sends $\lambda$ to $[x,S^{\lambda((2))}\otimes S^{\lambda((1,1))}]$.
Next, we consider the other extreme case when $d=2$. The wreath product $\Sigma_{m}\wr\Sigma_{2}$ is generally not a Coxeter group. We remark that irreducibles for the Hu algebras share the same index set $I^{H}_{m\wr 2}:=\{[\nu^{\prime},\nu^{\prime\prime}]~{}|~{}\nu^{\prime}\neq\nu^{\prime\prime}\in\Pi_{m}\}\sqcup\{[\nu,\nu]_{+},[\nu,\nu]_{-}~{}|~{}\nu\in\Pi_{m}\}$ with the irreducibles for $\mathbb{C}[\Sigma_{m}\wr\Sigma_{2}]$ (See [Hu02, LNX23]), where $[\nu^{\prime},\nu^{\prime\prime}]$ denotes the equivalence class in $\Pi_{m}^{2}$ under the obvious $\Sigma_{2}$-action.
Example 6.5.2.
The index set $I^{C}_{m\wr 2}=\{\lambda:\Pi_{m}\to\Pi~{}|~{}\sum_{\nu\vdash m}|\lambda(\nu)|=2\}$ is in bijection with $I^{H}_{m\wr 2}$ via
$\lambda\mapsto\begin{cases}[\nu,\nu]_{+}&\textup{if }\lambda(\nu)=(2);\\ [\nu,\nu]_{-}&\textup{if }\lambda(\nu)=(1,1);\\ [\nu^{\prime},\nu^{\prime\prime}]&\textup{if }\lambda(\nu^{\prime})=\lambda(\nu^{\prime\prime})=(1).\end{cases}$ (6.5.1)
On the other hand, any $x\in\mathcal{N}_{m}^{2}$ either has Jordan type $\nu^{2}$ or $\nu^{\prime}\nu^{\prime\prime}$ for some $\nu\in\Pi_{m}$ or $\nu^{\prime}\neq\nu^{\prime\prime}\in\Pi_{m}$. For the former case, $C(x)\cong\Sigma_{2}$ and hence $x$ corresponds to two distinct elements $[x,S^{(2)}]$ and $[x,S^{(1,1)}]$ in $I^{G}_{m\wr 2}$. For the latter case, $C(x)\cong\Sigma_{1}\times\Sigma_{1}$ and hence $x$ corresponds exactly one element $[x,S^{(1)}]$ in $I^{G}_{m\wr 2}$.
Finally, $\Psi$ sends $\lambda$ to $[x,S^{\lambda(\nu)}]$ if $|\lambda(\nu)|=2$ for some $\nu\vdash m$, $x\in\mathcal{N}_{m}^{2}$ with Jordan type $\nu^{2}$; and it sends $\lambda$ to $[x,S^{(1)}]$ if $\lambda(\nu^{\prime})=\lambda(\nu^{\prime\prime})=(1)$ for some $\nu^{\prime}\neq\nu^{\prime\prime}\vdash m$, $x\in\mathcal{N}_{m}^{2}$ with Jordan type $\nu^{\prime}\nu^{\prime\prime}$.